OpenStudy (anonymous):

Three pounds of cashews and four pounds of peanuts cost $25.75, while four pounds of cashews and two pounds of peanuts cost $26.00. Find the cost per pound of cashews.

OpenStudy (mertsj):

3c+4p=25.75 4c+2p=26

OpenStudy (mertsj):

Solve the system

OpenStudy (anonymous):

Call 1 pound of cashews c and one pound of peanuts p. Now make equations based on what you know about them. 3c+4p=25.75 Solve one of your equations for either c or p. Now, in the equation you didn't solve for one variable, replace that letter with what you came up with. Then solve for that letter.